OpenStudy (anonymous):

Good morning people! I am ready to give medals for help... y-intercept of f(x)=x^2-13x-4 divided by x^2+12-2

OpenStudy (anonymous):

good morning If!

OpenStudy (anonymous):

morning

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

You guys sure it's not like this: \[\Large f(x)=\frac{ x^2-13x-4 }{ x^2+12\color{red}x-2 }\] ?

OpenStudy (anonymous):

oops you are right terenz, it is, my oops

terenzreignz (terenzreignz):

Well, that's good. Carry on, you guys ^_^

OpenStudy (anonymous):

sorry if for the boo boo, it's the one above

OpenStudy (anonymous):

whoever helps will get a medal

OpenStudy (anonymous):

this is the problem f(x)=x^2−13x−4 divided by x^2+12x−2

OpenStudy (anonymous):

well my judgement is that when you've got the find y intercept you solve the equation for f(x) where x is 0 so: y=f(0)

OpenStudy (anonymous):

is that replace 0 where the x's are?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok wouldn't I have to factor first, it doesn't look like I can here

OpenStudy (anonymous):

you don't need any factorisation you just place 0 where x is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0, 2

OpenStudy (anonymous):

I hope you're saying y=2

OpenStudy (anonymous):

x=0, y=2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you got a well deserved medal if

OpenStudy (anonymous):

must remember to be exact and take my time when doing math...

OpenStudy (anonymous):

Thanks if! ;)

OpenStudy (anonymous):

You're welcome